Solve using the elimination method
-0.9x+0.8y=14.6
-0.3x+0.1y=3.7

Answer by palanisamy(496)

The given equations are,
-0.9x+0.8y=14.6 ...(1)
-0.3x+0.1y=3.7 ...(2)
(2)*3=> -0.9x+0.3y = 11.1 ..(3)
(1)-(3)=> 0.8y-0.3y = 14.6-11.1
0.5y = 3.5
y = 3.5/0.5
y = 7
(1)=> -0.9x+0.8*7 = 14.6
-0.9x = 14.6-5.6
-0.9x = 9
x = 9/(-0.9)
x = -10
Therefore the solution is x= -10 and y = 7